OpenStudy (anonymous):

How do you solve this? If a die is rolled five times, find the probability of getting at least one 6.

OpenStudy (anonymous):

are you saying 1/6th or just 1 6

OpenStudy (anonymous):

the probability of rolling a six would be 5/6

OpenStudy (anonymous):

since you are rolling it 5 times and there are 6 sides to a die

hartnn (hartnn):

the probability of getting atleast 1 6 = 1- probability of getting none of the 6 in any of the 5 rolls probability of not getting a 6 is 5/6 the probability of getting atleast 1 6 = 1-(5/6)^5 understood?

OpenStudy (anonymous):

one "6"

hartnn (hartnn):

u got how (5/6)^5 ?

OpenStudy (anonymous):

got it! thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (kropot72):

Using the binomial distribution the probability of throwing one 6 in five rolls is:\[P(one6)=\left(\begin{matrix}5 \\ 1\end{matrix}\right)\times \frac{1}{6}\times (\frac{5}{6})^{4}=(\frac{5}{6})^{5}\]